OpenStudy (mmae813):

Find the area of the region bounded by the graphs of y = x, y = −x + 4, and y = 0.

OpenStudy (irishboy123):

plot/draw it it looks like a triangle

OpenStudy (3mar):

Salam Your problem is plotted in the attached file. It is a triangle with its base on the x-axis. You can easily calculate the area of the traingle by the equation = \[\frac{ 1 }{ 2 }\times base \times height\] = \[\frac{ 1 }{ 2 }\times 4 \times 2 = 4\] It is interesting to use www.calorful.com for graghing, equation solving, and other cool thingss realted to Maths. Hope you got what you were looking for.

OpenStudy (3mar):

Thanks for the medal, IrishBoy123!